OpenStudy (anonymous):

find f(1) for f(x)= (e^ln1)/(x) ....is it f(1)=1?

OpenStudy (anonymous):

oops the ln in the equation should be lnx not ln1

OpenStudy (unklerhaukus):

like this? \[f(x)= \frac{e^{\ln x }}x\]\[f(1)= \frac{e^{\ln 1 }}1=\]

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

\[f(1)= \frac{e^{\ln 1 }}1=e^{\ln 1 }=e^{0}=\]

OpenStudy (anonymous):

ok thanks :)

OpenStudy (unklerhaukus):

yes =1